This one is stumping me.. I can't figure out how it comes out to be 16 and 32. Can someone show me the steps to get there? Thank you.

Jack is twice as old as Lacy. In 3 years the sum of the ages will be 54. How old are they now?

Answer by oberobic(2304)

J = Jack's current age
L = Lacey's current age
.
Jack is twice as old as Lacey.
J = 2L
.
In three years...
Jack will be J+3 years old, and
Lacey will be L+3 years old
.
...the sum of their ages will be 54.
(J+3)+(L+3) = 54
.
collect and simplify
.
J + L + 6 = 54
J + L = 48 (which is the sum of their current ages)
so,
J = 48-L
.
We know equals are equal, so we can substitute:
L = L, so
2L = 48-L
3L = 48
L = 16
.
Use that to find J.
J = 2L
J = 32
.
Check using the other facts given.
J+3 = 35
L+3 = 19
35 + 19 = 54
Correct.
.
Answer: Jack is 32, and Lacey is 16.
.
Done.